OpenStudy (afloridagirl):

The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan.

OpenStudy (afloridagirl):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^2+2x+8 is the same as 1x^2+2x+8 a = 1 b = 2 c = 8 you'll plug in the values of 'a' and 'b' into h = -b/(2*a) to get h = ???

OpenStudy (afloridagirl):

so it will look like 

OpenStudy (afloridagirl):

so 4

OpenStudy (afloridagirl):

is that wrong?

jimthompson5910 (jim_thompson5910):

c = 8 isn't used in h = -b/(2a)

jimthompson5910 (jim_thompson5910):

h = -b/(2a) h = -2/(2*1) h = ???

OpenStudy (afloridagirl):

oh im sorry so -1

jimthompson5910 (jim_thompson5910):

h = -1 is the x coordinate of the vertex plug x = -1 into y = x^2+2x+8 to get y = ???

OpenStudy (afloridagirl):

5 @jim_thompson5910

OpenStudy (afloridagirl):

im sorry i had too do something

jimthompson5910 (jim_thompson5910):

incorrect, try again

jimthompson5910 (jim_thompson5910):

x = -1 x^2 = (-1)^2 = (-1)*(-1) = +1

OpenStudy (afloridagirl):

oh hell, this is what i did